OpenStudy (anonymous):

c. If the area of a circle is 121π square units, what is its diameter? d. If the circumference of a circle is 20π units, what is its area?

OpenStudy (anonymous):

c) A=pi r^2 it will give you r=11, right? so, d=2r=22

OpenStudy (anonymous):

area of a circle is defined as \[ \pi r^2 \] and the diameter as 2r. Plugin the value and you can solve problem c. As for problem d, circumference is \[2 \pi r\], plugin the value and solve the equation. http://www.oojih.com

OpenStudy (anonymous):

hi jasywasy4 i had posted u the formula earlier, refr to them and work out. :)

OpenStudy (anonymous):

d) C=2pi r = 20pi it will give you r=... A=pi * r^2 =.... solve it

OpenStudy (anonymous):

c.\[Pi r^2=121\Pi hence r=11 hence d=22\] d, 2 pi r=20 pi r=10 hence area 100pi